You invented an air capacitor. Its parallel plates are separated by 2.00 mm. Each plate carries a charge of 5.50 nC. The magnitude of the electric field of the plates is 5.50 x 105 V/m. Find the capacitance of your invented air capacitor.



1
Expert's answer
2022-02-18T08:22:08-0500

Given:

"d=0.002\:\rm m"

"q=5.50*10^{-9}\:\rm C"

"E=5.50*10^5\:\rm V/m"


The capacitance

"C=\frac{q}{V}=\frac{q}{Ed}\\
=\frac{5.50*10^{-9}}{5.50*10^{5}*0.002}=5*10^{-12}\:\rm F"